OpenStudy (anonymous):

a(t)=p*e^rt how may years will it take for 50 dollar investment to grow to 1000 dollars compounded continuosly at a rate of 5%

OpenStudy (campbell_st):

so you have \[1000 = 50e^{0.05t}\] solve for t \[20 = e^{0.05t}\] take the base e log of both sides \[\ln (20) = 0.05t\] \[t = \frac{\ln(20)}{0.05}\] t = 59.9146 years

OpenStudy (anonymous):

@campbell_st so where did the 20 come from

OpenStudy (campbell_st):

20 = 1000/50

OpenStudy (anonymous):

@campbell_st oohhh thank you